OpenStudy (anonymous):

Sq root -1 over (2-4i) - (5+3i) =?

OpenStudy (anonymous):

\[\sqrt{\frac{ -1 }{ -3-7i }}=\sqrt{\frac{ 1*(3-7i) }{ (3+7i)(3-7i) }}=\sqrt{\frac{ 3-7i }{ 9+49 }}\] hope that helps you , n m sorry :(

OpenStudy (anonymous):

That's not even right.

OpenStudy (anonymous):

and if your question says \[\frac{ \sqrt{-1} }{ -3-7i }=\frac{ -i }{3+7i }*\frac{ 3-7i }{ 3-7i }=\frac{ 7-3i}{ 9+49 }\]

OpenStudy (anonymous):

this is the right one